# Multi-Margin Loss: Proposal and Application in Recommender Systems

### v1

@@ 3.2. Motivation for Multi-Margin Loss
For an efficient contrastive learning system, Weng et al.(Weng, 2021) identifies three key components: (i) Heavy Data Augmentation: Data augmentation techniques creates noisy versions of positive instances and are critical for learning generalizable embedding features. (ii) Large batch size: In order to cover a diverse and challenging negative samples, it is necessary to have larger batch sizes. (iii) Hard-negative mining: Hard negatives are the instances which are similar to the anchor item, i.e., user for recommender systems, but which are negative, i.e., user did not like that item. One solution for this purpose could be using large batch size or more negative samples, which implicitly introduce more hard negative samples to the system. However, it suffer from the usage of higher amount of memory as a side-effect. Another, more efficient and commonly used solution is to use margin $m$ for the computations. Using a margin $m$, it is possible to distinguish hard-negatives from the rest.
When examining the key components of contrastive learning, we recognized the importance of effectively utilizing negative samples while prioritizing the more challenging ones. We propose that using not just the most challenging instances but all non-trivial examples can lead to better model generalization, even with fewer negative samples or smaller batch sizes. Furthermore, we should be able to assign different weights based on the level of ”hardness” of negative samples, giving higher weight to the hardest negatives and relatively lower weights to easier ones. To address these considerations, we propose the Multi-Margin Loss, which is detailed in Section 3.3.
@@ 3.3. Multi-Margin Loss
4. Evaluation Results
This section presents the evaluation setup and the results.
@@ 4. Evaluation Results
4.1. Evaluation Setup
SimpleX (Mao et al., 2021a) showed that even with a simple neural-network architecture, a well-designed loss function can achieve good performance results. For comparisons, SimpleX (Mao et al., 2021a), which proposes CCL loss function, is used as the main comparison-partner. Also, recently proposed BSL (Wu et al., 2023) loss function, which outperforms CCL, is used as a comparison partner.
Datasets
Two datasets, which are commonly used in recommender systems, especially by recent neural-network-based CF models (Wang et al., 2019; Chen et al., 2020a; He et al., 2020; Mao et al., 2021a, b; Fan et al., 2023a; Zhou et al., 2023; Wang et al., 2024) are used: (i) Yelp2018 (Yelp, 2018) [^1] Yelp2018: https://huggingface.co/datasets/reczoo/Yelp18_m1 and (ii) Gowalla (Cho et al., 2011) [^2] Gowalla: https://huggingface.co/datasets/reczoo/Gowalla_m1. The statistics of the datasets are summarized in Table 1. Same as the SimpleX (Mao et al., 2021a), we used the datasets as is, without executing any preprocessing and/or filtering, such as filtering by popularity or cold-start.
Evaluation Metrics
Recall and NDCG are commong metrics used in recommender systems. For the performance comparison, we present the Recall@20 and NDCG@20 metrics, which are used by our comparison partners, CCL (Mao et al., 2021a) and BSL (Wu et al., 2023), as well.
Table 2. Performance comparison of MML and CCL on SimpleX and MF architectures with and without early-stopping (Number of negative samples = 100)
Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets.
@@ 4.1. Evaluation Setup
For the implementation of MML, the code provided by SimpleX (Mao et al., 2021a) is expanded and shared online [^3] MML (Proposed) code: TO-BE-SHARED. The experiments involving CCL utilized the code from the SimpleX (Mao et al., 2021a) [^4] SimpleX (CCL) (Mao et al., 2021a) code: https://reczoo.github.io/SimpleX. For experiments related to BSL, we used the code provided in the BSL (Wu et al., 2023) paper [^5] BSL (Wu et al., 2023) code: https://github.com/junkangwu/BSL. In the experiments, we used the same parameters as the SimpleX (Mao et al., 2021a) paper. For both the Yelp and Gowalla datasets, we used the Adam optimizer with $L_{2}$ regularization. The learning rate for both datasets was set to 1e-4, with an embedding dimension of 64 and 100 epochs. The batch size for the Yelp dataset was 512, with a regularizer of 1e-09, while for Gowalla, the batch size was 256, with a regularizer of 1e-08. We explored various numbers of negative samples, but results are presented using 100 negatives, unless otherwise specified.
@@
The MML-specific parameters were determined by exploring the parameter space. For the Yelp dataset, the margin values were set as [0.6, 0.7, 0.8, 0.9] with corresponding weights of [0.05, 0.15, 0.2, 0.6]. The weight for positive instances, $w_{p}$, was set to 1, and for negative instances, $w_{n}$, it was set to 400. For the Gowalla dataset, the margin values were [0.7, 0.8, 0.9], with weights of [0.10, 0.15, 0.75]. The weight for positive instances remained 1, and for negative instances, it was set to 700. For the CCL experiments, we used the original code and configuration, with the margin kept at 0.9. Unlike the original setup, we did not use early-stopping, as it was shown to produce better results for our datasets. Additionally, we did not apply learning-rate reduction to align with the BSL implementation, which doesn’t support this feature. In the BSL experiments, we retained the original code and configuration. The temperatures $\tau_{1}and\tau_{2}$ set as 0.11 and 1.10 for the Yelp dataset, and 0.08 and 1.22 for the Gowalla dataset. The original code specified 1000 epochs, but for consistency with the other experiments, we set the number of epochs to 100.
(a) Yelp dataset
(b) Gowalla dataset
Figure 3. Performance of MF model under recent loss functions while using various number of negative-samples.
Increasing number of negative-samples improves the performance for all loss functions. The proposed loss function, MML, outperforms existing baseline loss functions when the number of negative samples is 100 or fewer, suggesting it is more efficient in resource-constrained environments.
4.2. Evaluation Results
Impact of model architecture and early-stopping
The table reveals that for both the CCL and MML loss functions, and for both the SimpleX and MF architectures, omitting early-stopping leads to better performance. Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets. For the Yelp dataset, without early-stopping, MML outperforms CCL by 3.98% in recall and 4.14% in NDCG on SimpleX and by 4.31% in recall and 5.26% in NDCG on MF. On the Gowalla dataset, the gains with MML are 7.46% in recall and 7.05% in NDCG on SimpleX, while on MF, the improvement is 1.85% in recall and 0.63% in NDCG. We attribute the relatively lower performance gain on Gowalla dataset to the sparsity of the Gowalla dataset. In this dataset, MML can effectively learn from the available negative instances by considering all non-trivial negatives. However, due to the sparsity of the dataset, MML may struggle to learn sufficiently from the positive instances.
Impact of number of negative samples
Table 3. Performance of MF model with different loss functions (Number of negative samples = 100)
@@ 4.2. Evaluation Results
MML performs equally well in recall and NDCG on the Yelp dataset and achieves the best results in recall on the Gowalla dataset.

### v2

@@ 3.2. Motivation for Multi-Margin Loss
For an efficient contrastive learning system, Weng et al.(Weng, 2021) identify three key components: (i) Heavy Data Augmentation: Data augmentation techniques create noisy versions of positive instances and are critical for learning generalizable embedding features. (ii) Large Batch Size: In order to cover a diverse and challenging negative samples, it is necessary to have larger batch sizes. (iii) Hard-Negative Mining: Hard negatives are the instances which are similar to the anchor item, i.e., user for recommender systems, but which are negative, i.e., user did not like that item. One solution for incorporating hard negatives is to use large batch sizes or more negative samples, which implicitly introduce more hard negative samples to the system. However, this approach suffers from higher memory usage as a side effect. A more efficient and commonly used solution is to use a margin $m$ for the computations. With a margin $m$, it is possible to distinguish hard negatives from the rest.
When examining the key components of contrastive learning, we recognized the importance of effectively utilizing negative samples while prioritizing the more challenging ones. Instead of relying only on larger batch sizes to introduce more hard negatives, which can be memory-intensive, we propose that using not just the most challenging instances but all non-trivial examples can lead to better model generalization, even with fewer negative samples or smaller batch sizes. We propose assigning different weights based on the ”hardness” of negative samples, with higher weights given to the hardest negatives and relatively lower weights to easier ones. This prioritization helps the model to distinguish between different levels of ”hardness” among negative samples, effectively using a smaller number of negative samples. Furthermore, considering all non-trivial negative examples enhances generalization, ensuring the model learns from a diverse set of negative samples.
Training a recommender model using contrastive learning helps distinguish between similar and dissimilar pairs and learn representations that capture meaningful relationships between users and items. Better representation of users and items can improve recommendation performance by generalizing well even with limited interaction data, understanding a broader spectrum of item similarities and dissimilarities, and ensuring that the negative pairs represent diverse and unbiased samples. To address these considerations, we propose the Multi-Margin Loss, which is detailed in Section 3.3.
@@ 3.3. Multi-Margin Loss
Table 2. Performance comparison of MML and CCL on SimpleX and MF architectures with and without early-stopping (Number of negative samples = 100). Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets.
@@
Table 3. Training-Runtime Comparison on MF model (Without early-stopping, 100 epochs). MML increases training time per epoch by at most 2 seconds on average.
#Negative samples Loss Yelp18 Gowalla
Time per epoch Ratio Total time Time per epoch Ratio Total time
800 CCL (2021) 77.5 secs - $\sim$129 mins 66 secs - 110 mins
MML (2024) - Proposed 78.5 secs +1.3% $\sim$130 mins 64 secs -3.0% $\sim$107 mins
100 CCL (2021) 35 secs - $\sim$58 mins 37 secs - $\sim$62 mins
MML (2024) - Proposed 35 secs 0.0% $\sim$58 mins 37.5 secs +1.4% $\sim$63 mins
10 CCL (2021) 31 secs - $\sim$51 mins 32 secs - $\sim$53 mins
MML (2024) - Proposed 33 secs +6.5% 55 mins 33.5 secs +4.7% $\sim$56 mins
4. Evaluation Results
This section presents the evaluation setup and the results.
4.1. Evaluation Setup
SimpleX (Mao et al., 2021a) demonstrated that even with a simple neural-network architecture, a well-designed loss function can achieve good performance results. For comparisons, SimpleX (Mao et al., 2021a), which proposes the CCL loss function, is used as the main benchmark. In addition to the CCL loss, five more loss functions are used for comparisons: (i) four classical loss functions: BPR loss, Pairwise Hinge loss, Softmax Cross-Entropy loss, and Mean Square Error loss; (ii) two recently proposed contrastive loss functions: CCL (Mao et al., 2021a) and BSL (Wu et al., 2023).
Datasets and Evaluation Metrics
Two datasets, which are commonly used in recommender systems, especially by recent neural-network-based CF models (Wang et al., 2019; Chen et al., 2020a; He et al., 2020; Mao et al., 2021a, b; Fan et al., 2023a; Zhou et al., 2023; Wang et al., 2024) are used: (i) Yelp2018 (Yelp, 2018) [^1] Yelp2018: https://huggingface.co/datasets/reczoo/Yelp18_m1 and (ii) Gowalla (Cho et al., 2011) [^2] Gowalla: https://huggingface.co/datasets/reczoo/Gowalla_m1. These datasets have already been split, pre-processed, and made publicly available. The statistics of the datasets are summarized in Table 1. Same as the SimpleX (Mao et al., 2021a), we used the datasets as is, without executing any preprocessing and/or filtering, such as filtering by popularity or cold-start. We consider all items not interacted with by a user as candidates and report the average results across all users.
In order to evaluate the performance of recommendations, we use commonly accepted metrics like Recall and NDCG. In this work, we specifically present the Recall@20 and NDCG@20 metrics for comparing performance, which are also used by our comparison partners, CCL (Mao et al., 2021a) and BSL (Wu et al., 2023).
(a) Yelp dataset
(b) Gowalla dataset
Figure 3. Performance of MF model under recent loss functions while using various number of negative-samples. Increasing number of negative-samples improves the performance for all loss functions. The proposed loss function, MML, outperforms existing baseline loss functions when the number of negative samples is 100 or fewer, suggesting it is more efficient in resource-constrained environments.
@@ 4.1. Evaluation Setup
For the implementation of MML, the code provided by SimpleX (Mao et al., 2021a) is expanded and shared online [^3] MML (Proposed) code: TO-BE-SHARED. The experiments involving CCL utilized the code from the SimpleX (Mao et al., 2021a) [^4] SimpleX (CCL) (Mao et al., 2021a) code: https://reczoo.github.io/SimpleX. For experiments related to BSL, we used the code provided in the BSL (Wu et al., 2023) paper [^5] BSL (Wu et al., 2023) code: https://github.com/junkangwu/BSL. Although the original code implementations are utilized, all the experiments have been re-run and the results are presented in the following sections. In the experiments, we used the same parameters as the SimpleX (Mao et al., 2021a) paper. For both the Yelp and Gowalla datasets, we used the Adam optimizer with $L_{2}$ regularization. The learning rate for both datasets was set to 1e-4, with an embedding dimension of 64 and 100 epochs. The batch size for the Yelp dataset was 512, with a regularizer of 1e-09, while for Gowalla, the batch size was 256, with a regularizer of 1e-08. We explored various numbers of negative samples, but results are presented using 100 negatives, unless otherwise specified.
@@
The MML-specific parameters were determined through grid search across the parameter space. We explored $w_{p}$:$w_{n}$ ranging from [1:150] to [1:450] in increments of 50 for Yelp and from [1:500] to [1:1000] in increments of 100 for Gowalla. Initial margin values started at [0.9] and were expanded until [0.5, 0.6, 0.7, 0.8, 0.9] for both datasets. Margin weights were adjusted with priority given to the highest margin and reductions made proportionally. As a result, for Yelp, margins were set to [0.6, 0.7, 0.8, 0.9] with corresponding weights of [0.05, 0.15, 0.2, 0.6], and $w_{p}$:$w_{n}$ was set to [1:400]. For Gowalla, margins were set to [0.7, 0.8, 0.9] with weights of [0.10, 0.15, 0.75], and $w_{p}$:$w_{n}$ was set to [1:700]. We want to use AutoML tools for searching the right parameters as a future direction. For the CCL experiments, we used the original code and configuration, with the margin kept at 0.9. Unlike the original setup, we did not use early-stopping, as it was shown to produce better results for our datasets. Additionally, we did not apply learning-rate reduction to align with the BSL implementation, which doesn’t support this feature. In the BSL experiments, we retained the original code and configuration. The temperatures $\tau_{1}and\tau_{2}$ set as 0.11 and 1.10 for the Yelp dataset, and 0.08 and 1.22 for the Gowalla dataset. The original code specified 1000 epochs, but for consistency with the other experiments, we set the number of epochs to 100.
@@
Table 4. Performance of MF model with different loss functions (Number of negative samples = 100). MML performs equally well in recall and NDCG on the Yelp dataset and achieves the best results in Recall on the Gowalla dataset.
@@
4.2. Evaluation Results
Impact of model architecture and early-stopping
The table demonstrates that in our experiment omitting early-stopping results in better performance for both the CCL and MML loss functions, as well as for both the SimpleX and MF architectures. Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets. For the Yelp dataset, without early-stopping, MML outperforms CCL by 3.98% in recall and 4.14% in NDCG on SimpleX and by 4.31% in recall and 5.26% in NDCG on MF. On the Gowalla dataset, the gains with MML are 7.46% in recall and 7.05% in NDCG on SimpleX, while on MF, the improvement is 1.85% in recall and 0.63% in NDCG. We attribute the relatively lower performance gain on Gowalla dataset to the sparsity of the Gowalla dataset. In this dataset, MML can effectively learn from the available negative instances by considering all non-trivial negatives. However, due to the sparsity of the dataset, MML may struggle to learn sufficiently from the positive instances.
Training Runtime Comparison
We compare the training runtime of the CCL and MML loss functions using MF architecture in Table 3. The table shows the average time spent per epoch, change ratio and total training time on the Yelp and Gowalla datasets using varying numbers of negative samples. All the experiments use the same setup described in the Implementation Details in Section 4.1.
The Table 3 shows that for both CCL and MML loss functions the time spent per epoch increases as the number of negative samples increases. Training a model with 800 negative samples takes about 2 hours, but it takes around 1 hour with 100 negative samples. Comparing CCL and MML shows that using multiple margins has a relatively small impact, with the training time per epoch increasing by no more than 2 seconds on average. This small increase in training time, while using negative samples more efficiently, makes MML a promising loss function for recommender systems.
Impact of number of negative samples